40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 13 14:12:03 2022, Yehonathan, Update, BHD, Trying LO phase locking again Screenshot_2022-09-13_17-41-33.png
    Reply  Thu Sep 15 21:12:53 2022, Paco, Update, BHD, LO phase "dc" control lophase_oltf.pdflophase_noise_spectra.pdf
       Reply  Tue Sep 27 10:50:11 2022, Paco, Update, BHD, calibrated LO phase noise lo_phase_asd.pdf
          Reply  Wed Sep 28 16:37:26 2022, Paco, Update, BHD, calibrated LO phase noise; update lo_phase_asd.pdf
             Reply  Wed Sep 28 21:54:08 2022, Paco, Update, BHD, calibrated LO phase noise; update dcpd_phase_asd.pdflo_phase_asd.pdf
                Reply  Fri Sep 30 20:18:55 2022, Paco, Update, BHD, LO phase noise with different actuation points BHDC_asd_actuation_point.png
                   Reply  Mon Oct 3 15:19:05 2022, Paco, Update, BHD, LO phase noise and control after violin mode filters filters_c1sus_lo1_lsc.pngBHDC_asd_act.pngboosted_lo_phase_control.png
Message ID: 17161     Entry time: Wed Sep 28 16:37:26 2022     In reply to: 17160     Reply to this: 17163
Author: Paco 
Type: Update 
Category: BHD 
Subject: calibrated LO phase noise; update 

[yuta, paco]

Update; the high frequency ( > 100 Hz) drop is of course not real and comes from a 4th order LP filter in the HPC demod I filter which I haven't accounted for. Furthermore, we have gone through the calibration factors and corrected a factor of 2 in the optical gain. Then, I also added the CLTF to show in loop and out of loop error respectively. The updated plot, though not final, is in Attachment #1.

Attachment 1: lo_phase_asd.pdf  204 kB  Uploaded Wed Sep 28 17:39:37 2022  | Hide | Hide all
lo_phase_asd.pdf
ELOG V3.1.3-